### v1

# Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling

Abstract
Constitutive modeling is crucial for engineering design and simulations to accurately describe material behavior. However, traditional phenomenological models often struggle to capture the complexities of real materials under varying stress conditions due to their fixed forms and limited parameters. While recent advances in deep learning have addressed some limitations of classical models, purely data-driven methods tend to require large datasets, lack interpretability, and struggle to generalize beyond their training data. To tackle these issues, we introduce "Fusion-based Constitutive model (FuCe): Towards model-data augmentation in constitutive modelling". This approach combines established phenomenological models with an ICNN architecture, designed to train on the limited and noisy force-displacement data typically available in practical applications. The hybrid model inherently adheres to necessary constitutive conditions. During inference, Monte Carlo dropout is employed to generate Bayesian predictions, providing mean values and confidence intervals that quantify uncertainty. We demonstrate the model’s effectiveness by learning two isotropic constitutive models and one anisotropic model with a single fiber direction, across six different stress states. The framework’s applicability is also showcased in finite element simulations across three geometries of varying complexities. Our results highlight the framework’s superior extrapolation capabilities, even when trained on limited and noisy data, delivering accurate and physically meaningful predictions across all numerical examples.
Keywords Model-Data Fusion, Hybrid Constitutive model, ICNN, Hyper-elasticity, Approximate Bayesian, Physics enhanced.
1 Introduction
In simulations of physical systems, constitutive modeling is essential for characterizing the mechanical behavior of materials. The accuracy of these simulations heavily depends on choosing the right constitutive model [1]. While some models are grounded in physics, most are primarily phenomenological. These models work well for small deformations in materials with linear behavior. However, when dealing with large deformations in materials like rubbers, polymers, foams, and biological tissues –which exhibit nonlinear and complex behaviors, developing accurate empirical or phenomenological models becomes significantly more difficult. Even with careful formulation and calibration, these models are limited by their fixed forms and the small number of parameters, restricting their ability to accurately represent a wide range of deformation levels and stress states. Machine learning based data-driven approach [2] including methods like Gaussian process regression [3, 4] or neural networks (NNs) [5, 6] presents a reasonable alternatives to overcome the time consuming task of formulating classical constitutive models and to improve the limited expressibility associated with these phenomenological models. The concept of using NNs in constitutive modeling was first presented in early 1990s [7]. The fundamental concept of machine-learning-based material modeling involves choosing a highly expressive model approach, characterized by a large number of learnable parameters, and then tuning these parameters using the available data. However, in this early phase, black-box approaches were mostly used, i.e., networks that do not take into account any physical principles and therefore only performs satisfactorily in the training regime. However, these models offer poorly generalization beyond the training data regime. Moreover, data-driven constitutive modeling requires a large number of strain-stress pairs. This presents two significant challenges, especially when relying on experimental data. First, it is impractical to explore the entire high-dimensional strain-stress space using standard mechanical tests, such as uni-/bi-axial tensile or bending tests. Second, complete tensorial constitutive models are difficult to develop because stress tensors cannot be directly measured; force measurements are only boundary-averaged projections of stress tensors. While multiscale simulations can generate extensive and detailed stress-strain data [8, 9], the computational cost for complex material systems is currently prohibitive. Therefore, it is crucial to learn material behavior directly from the force-displacement data that is practically available through mechanical testing.
Modern material model calibration primarily relies on full-field displacement data, such as that obtained via digital image correlation (DIC), combined with applied force data from load cells [10]. This calibration is typically performed using the finite element model updating (FEMU) [11] or virtual fields method (VFM) [12]. VFM directly solves for unknown material parameters by satisfying the weak form of momentum balance with measured displacement data, while FEMU iteratively optimizes model parameters until the simulated and measured displacement fields match. Recent efforts have demonstrated the estimation of stress fields using full-field displacement data in a model-free data-driven context [13, 14, 15]. In the model-based domain, physics-informed neural networks (PINNs) and their variants have shown promising results by first learning the forward solution to the mechanical boundary value problem and then estimating unknown parameters via gradient-based optimization [16, 17, 18, 19]. However, all these techniques, including FEMU, VFM, and PINNs, are limited to predefined constitutive models with a few unknown parameters. Such models are specific to particular materials and stress states and cannot be generalized beyond the calibration data.
The limitations of the previously described frameworks can be mitigated by adopting a hybrid framework through data-model fusion [20]. These models integrate phenomenological and data-driven approaches. Various hybrid strategies exist for combining data-driven models with physics-based models to solve partial differential equations (PDEs). For instance, Chakraborty [21] introduced a transfer learning-based multi-fidelity physics-informed deep learning framework, which demonstrated efficiency but lacked interpretability. Our recent works [22, 23], also proposed an alternative framework that augments known physics with deep learning models for solving partial differential equations (PDEs) and stochastic differential equations (SDEs). These hybrid frameworks are found to be accurate, more interpretable and generalizes well to unseen environments.
In this research work, our objective is to develop a hybrid framework for constitutive modeling utilising existing phenomenological models along with the deep learning frameworks, using only the global reaction force and full-field displacement data. This approach is crucial due to the difficulty in availing stress strain data, practical problem of data scarcity, the difficulty in formulating accurate phenomenological models, and the critical need for generalization when a material’s behavior changes under varying external conditions. The salient features of the proposed FuCe approach are as follows:
  • Model-data fusion: The proposed framework utilizes classical phenomenological constitutive models alongside deep learning models, which augment and correct the phenomenological models to allign with real material data.
  • Satisfies all necessary constitutive conditions: By employing an ICNN neural network with appropriate correction terms, the model satisfies all thermodynamic and physical feasibility conditions.
  • Utilizes practically feasible force-displacement noisy data: The framework uses practically available full-field displacement and global reaction force data for training purposes, accounting for noise in the data.
  • Physics-informed loss: The training process is guided by a physics-informed loss function, compensating for the lack of direct stress/strain energy data.
  • Approximate Bayesian inference: To quantify uncertainties in model predictions [24], a Bayesian approximation method called Monte Carlo Dropout [25] is used during the inference stage.
Overall, the proposed framework is a novel hybrid model of its kind, that utilises deep learning models to learn the discrepancy between the known phenomenological model and the actual material behaviour utilizing the practically available full-field displacement and global reaction force data.
The remainder of the paper is organized as follows. The background of constitutive models specifically hyperelasticity is formally described in Section 2. Detailed description on the problem statement is also presented at the end of Section 2. The proposed approach along with the algorithm is described in Section 3. Results and discussion of the numerical experiments that are carried out, are presented in Section 4. Finally, Section 5 presents the concluding remarks.
2 Background and problem statement
2.1 Stress-Deformation measures and Hyper-elasticity
Consider a solid body, having a reference configuration $\mathbf{X}\subset\mathbb{R}^{3}$ at initial time $t_{0}\in\mathbb{R}$. After deformation, at time $t\in[t_{0},\infty)$, the body attains its current configuration $\mathbf{x}\subset\mathbb{R}^{3}$. This motion is described by a bijective mapping $\bm{\zeta}:\mathbf{X}\times[t_{0},\infty)\rightarrow\mathbf{x}$, during which a typical material particle in the reference configuration $\mathbf{X}$ undergoes a displacement $\mathbf{u}(\mathbf{X},t)$ to arrive at their positions in the current configuration $\mathbf{x}$, such that:
$\mathbf{x}=\bm{\zeta}(\mathbf{X},t)=\mathbf{X}+\mathbf{u}(\mathbf{X},t)\quad\text{with}\quad\mathbf{X}=\bm{\zeta}(\mathbf{X},t_{0})$ (1)
With regards to the aforementioned deformation, we define some deformation measures. At first, the gradient of the function $\bm{\zeta}(\mathbf{X},t)$ is known as the deformation gradient $\bm{F}\in\mathcal{GL}^{+}(3)$ is defined as:
$\bm{F}=\nabla_{\mathbf{X}}\bm{\zeta}=\mathbb{I}+\nabla\mathbf{u}$ (2)
and its determinant as $J=\frac{dV}{dV_{0}}=\det\bm{F}\in\mathbb{R}^{+}$, where $\mathcal{GL}^{+}(3)$ denotes the set of all invertible second-order tensors with positive determinant. Another deformation measure, ensuring the absence of rigid body motions, is represented by the positive definite right Cauchy–Green deformation tensor:
$\mathbf{C}=\bm{F}^{T}\bm{F}$ (3)
Lets now define some stress measures, starting with the true stress or Cauchy stress tensor $\bm{\sigma}$. It is a second-order symmetric tensor that describes the stress vector (a.k.a traction vector) $\mathbf{t}$ resulting from an infinitesimal force $d\mathbf{f}$ applied on an infinitesimal surface $d\mathbf{A}$ with normal $\mathbf{n}$ in the current configuration:
$\mathbf{t}=\lim_{d\mathbf{A}\to 0}\frac{d\mathbf{f}}{d\mathbf{A}}=\bm{\sigma}\mathbf{n}$ (4)
The infinitesimal force $d\mathbf{T}$ can be expressed either by taking area $d\mathbf{S}$ and its normal $\mathbf{n}$ considering the current configuration or by taking area $d\mathbf{S}_{0}$ and its normal $\mathbf{n}_{0}$ of the original (initial) configuration. This can be expressed as:
$d\mathbf{f}=\bm{\sigma}\mathbf{n}\,d\mathbf{A}=\bm{P}\mathbf{n}_{0}\,d\mathbf{A}_{0}$ (5)
where $\bm{P}$ is known as the (first) Piola-Kirchhoff stress tensor. The first Piola-Kirchhoff stress tensor $\mathbf{P}$ cam also be related to the Cauchy stress tensor $\bm{\sigma}$ through the deformation gradient $\bm{F}$ and the Jacobian $J$:
$\mathbf{P}=J\sigma\bm{F}^{-T}$ (6)
Another measure of stress can be expressed by defining the undeformed force $d\mathbf{\tilde{f}}=\bm{F}^{-1}d\mathbf{f}$, which then can be defined as:
$d\mathbf{\tilde{f}}=\bm{S}\mathbf{n_{0}}\,d\mathbf{A_{0}}$ (7)
where $\bm{S}$ denotes the second Piola-Kirchhoff stress tensor. It is a symmetric stress tensor that can be written as:
$\mathbf{S}=J\bm{F}^{-1}\bm{\sigma}\bm{F}^{-T}=\bm{F}^{-1}\bm{P},$ (8)
with $\rho_{0}$ and $\rho$ representing the mass densities of the reference and current configuration, respectively, accounting for mass conservation, $\rho_{0}=J\rho$ yields the balance of linear momentum with regard to the reference configuration:
$\nabla_{\mathbf{X}}\cdot\mathbf{P}+\rho_{0}\mathbf{f}=0.$ (9)
The preceding equation does not include inertia factors, and the mass-specific force density is indicated by $\mathbf{f}$. Furthermore, the balance of angular momentum corresponds to:
$\mathbf{P}\cdot\mathbf{F}^{T}=\mathbf{F}\cdot\mathbf{P}^{T},$ (10)
when taking into account the balances of mass and linear momentum.
2.2 General criteria for hyper-elasticity
In elasticity, the goal of constitutive modeling is to establish a relationship between the strain at a material location and the corresponding stress. However, in hyperelastic constitutive modeling, instead of a direct mapping, the stress-strain relationship is derived through a potential $\psi\colon\mathcal{GL}^{+}(3)\to\mathbb{R}$, $\bm{F}\to\psi(\bm{F})$, which corresponds to the strain energy density stored in the body. This relationship can be expressed as:
$\bm{P}(\bm{F})=\frac{\partial\psi(\bm{F})}{\partial\bm{F}}\quad\text{and}\quad\mathbb{C}(\bm{F})=\frac{\partial\bm{P}(\bm{F})}{\partial\bm{F}},\quad\forall\,\bm{F}\in\mathcal{GL}^{+}(3)$ (11)
where $\mathbb{C}(\bm{F})$ denotes the (incremental) tangent modulus. Thus, the goal of hyperelastic constitutive modeling is to learn this strain energy potential $\psi(\bm{F})$. However, for any learned material model $\psi(\bm{F})$ to be mathematically plausible, it must also meet the following thermodynamic and physical requirements:
  • Thermodynamic consistency: The stress tensor is thermodynamically consistent by construction since its definition as a gradient field in Eq. (11) entails energy conservation and path independence [26, 27, 28].
  • Stress tensor symmetry: The construction of $\psi(\bm{F})$ must guarantee consistency with the angular momentum balance. This requirement can be represented as follows by applying Eq. (10) and (11):
    $\frac{\partial\psi}{\partial\bm{F}}\cdot\bm{F}^{T}=\bm{F}\cdot\frac{\partial\psi}{\partial\bm{F}^{T}}$ (12)
    This leads to the necessity for the symmetry of the stress tensors $\bm{\sigma}$ and $\bm{S}$, using the stress transformations presented in Section 2.1 for eg:
    $\bm{S}=\bm{F}^{-1}\cdot\frac{\partial\psi}{\partial\bm{F}}\overset{!}{=}\frac{\partial\psi}{\partial\bm{F}^{T}}\cdot\bm{F}^{-T}=\bm{S}^{T}$ (13)
  • Objectivity: The constitutive model must be independent of the choice of material frame i.e:
    $\psi(\bm{R}\bm{F})=\psi(\bm{F}),\quad\forall\bm{F}\in\mathcal{GL}^{+}(3),\bm{R}\in\text{SO}(3),$ (14)
    where $\text{SO}(3)$ denotes the 3D rotation group.
  • Normalisation condition for stress and energy: In the absence of any deformation, the stress and the strain energy must be zero i.e.,
    $\psi(\bm{F}=\bm{I})=\bm{P}(\bm{F}=\bm{I})=\mathbf{0}$ (15)
  • Polyconvexity The concept of polyconvexity in the calculus of variations is generalisation of the concept of quasicovexity. The material must be quasi-convex in order for it to be stable [29, 30]. Since, imposing quasi-convexity is not numerically tractable, this is relaxed by enforcing polyconvexity [31, 30]. $\psi(\bm{F})$ is considered polyconvex if there exists a convex function $\mathcal{P}$ such that
    $\psi(\bm{F})=\mathcal{P}(\bm{F},\text{Cof}\bm{F},\det\bm{F})$ (16)
2.3 Classical Phenomenological models
In the study of hyperelastic materials, strain energy density functions are often defined using various deformation invariants such as the principal invariants of the right Cauchy-Green deformation tensor $\bm{C}$: $I_{1}=\mathrm{tr}(\bm{C})$, $I_{2}=\frac{1}{2}[\mathrm{tr}(\bm{C})^{2}-\mathrm{tr}(\bm{C}^{2})]$, and $I_{3}=\det(\bm{C})$. While most models were originally formulated with the assumption of incompressibility, it is crucial for constitutive models to account for compressibility, therefore we have assumed a deviatoric-volumetric split. The volumetric invariant $J$ is given by the determinant of the deformation gradient $\bm{F}$, specifically $J=\det(\bm{F})=I_{1}^{1/2}$. In isotropic materials, the deviatoric invariants are $\tilde{I}_{1}=J^{-2/3}I_{1}$ and $\tilde{I}_{2}=J^{-4/3}I_{2}$ and for anisotropic materials the deviatoric invariants are $\tilde{I}_{\alpha_{i}}=J^{-2/3}(\mathbf{a}^{(i)}\cdot\bm{C}\mathbf{a}^{(i)})$, where $\mathbf{a}^{(i)}$ denotes the $i^{th}$ fiber direction [32, 33]. In case of unknown fiber orientations $\{\alpha_{i}:i=1,2,\ldots\}$, the fiber directions $\mathbf{a}^{(i)}=(\cos\alpha_{i},\sin\alpha_{i},0)^{T}$ are treated as trainable parameters.
The invariants are further shifted by appropriate scalar values, and in some cases squared, to ensure that their values and the respective derivatives with respect to $\mathbf{F}$ are zero in case of no deformation ($\mathbf{F}=\mathbf{I}$). In this work, we utilize specific forms of strain energy density functions from classical models for data generation and training purposes, as detailed below:
  1. 1.
    Neo-Hookean (NH) Model [34]:
    $\psi(\mathbf{F})=0.5(\tilde{I}_{1}-3)+1.5(J-1)^{2}$ (17)
  2. 2.
    Isihara (IH) Model [35] :
    $\psi(\mathbf{F})=0.5(\tilde{I}_{1}-3)+(\tilde{I}_{2}-3)+(\tilde{I}_{1}-3)^{2}+1.5(J-1)^{2}$ (18)
  3. 3.
    Arruda–Boyce (AB) Model [36]:
    $\psi(\mathbf{F})=2.5\sqrt{N_{c}}\left[\beta_{c}\lambda_{c}-\sqrt{N_{c}}\log\left(\frac{\sinh\beta_{c}}{\beta_{c}}\right)\right]-c_{AB}+1.5(J-1)^{2}$ (19)
    where $\lambda_{c}=\left(\frac{\tilde{I}_{1}}{3}\right)^{1/2}$, $\beta=\mathcal{L}^{-1}\left(\frac{\lambda_{c}}{\sqrt{N_{c}}}\right)$, and $\mathcal{L}^{-1}$ denotes the inverse Langevin function. The constants are set to $N_{c}=28$ (denoting the number of polymeric chain segments) and $c_{AB}\approx 3.7910$, the latter used to offset the energy density to zero at $\mathbf{F}=\mathbf{I}$ (since the Arruda–Boyce feature does not itself vanish for zero deformation).
  4. 4.
    Anisotropic Model with one fiber family at $\alpha_{1}=45^{\circ}$ orientation (AI45):
    $\psi(\mathbf{F})=0.5(\tilde{I}_{1}-3)+0.75(J-1)^{2}+0.5(\tilde{I}_{\alpha_{1}}-1)^{2}$ (20)
These models struggle to capture the complex behaviors of real materials due to their fixed form and restrictive functional dependence on invariants. The difficulty further escalates when trying to meet all the physical constraints outlined in Section 2.2. Additionally, each model has specific limitations under varying deformation levels and loading conditions. For example, the Neo-Hookean model is advantageous due to its simplicity and computational efficiency; however, it requires stringent conditions to achieve accurate results. The Mooney-Rivlin model performs well for small to medium strains, while the Gent model excels in predicting large strain loadings (up to 300%). The Isihara model is accurate for moderate stretch values but poorly estimates behavior under biaxial loading. The Ogden model is highly suitable for large deformation behavior; however, its drawback is the need for different sets of material parameters for each deformation mode.
Addressing these limitations can be achieved through the use of machine learning models. Traditional models often require extensive stress-strain data, which is not always practical. It is crucial to train the model using realistically measurable full-field displacement and global reaction force data. End-to-end training with force-displacement data is challenging and necessitates a large dataset. Furthermore, these models struggle to generalize beyond the training regime. A logical step forward is to integrate existing phenomenological models with the data driven models. This hybrid model can train effectively with less data and generalize well for large strains, overcoming the limitations of the traditional methods.
2.4 Problem statement
Consider a 2D plate in a two-dimensional domain $\Omega$ with boundary $\partial\Omega$ is subjected to quasi-static loading. We apply Dirichlet boundary conditions on $\partial\Omega_{u}\subseteq\partial\Omega$ and Neumann boundary conditions on $\partial\Omega_{t}=\partial\Omega\setminus\partial\Omega_{u}$. The tests are conducted using displacement controlled setting, ensuring homogeneous Neumann boundary conditions and matching the forces applied by the testing equipment with the reaction forces at the fixed boundary. The available data includes displacement measurements $\{\bm{u}^{t}_{r}:r=1,...,n_{d};t=1,...,n_{l}\}$ at $n_{d}$ points and net reaction force measurements $\{\bm{R}^{t}_{\beta}:\beta=1,...,n_{\beta};t=1,...,n_{l}\}$ on some boundary segments, collected over $n_{l}$ load steps. The net reaction forces on segments of $\partial\Omega_{u}$ (e.g., each loaded side of the specimen) are assumed to be known, simulating measurements from load cells. With this setup, the overall objective is to discover the underlying constitutive law of the unknown material. We have prior knowledge that the material exhibits hyperelastic behavior, and a potential constitutive model, $\psi_{\text{known}}(\bm{F})$ is also available. However, as is often the case, $\psi_{\text{known}}(\bm{F})$ is approximate, and the true constitutive relation takes the following form
$\psi(\bm{F})=\psi_{\text{known}}(\bm{F})+\psi_{\text{missing}}(\bm{F}).$ (21)
cannot rely solely on phenomenological models. Since the constitutive model $\psi_{\text{known}}(F)$ is not accurate, the obtained stress field will also be inaccurate, leading to an erroneous force field solution. With this setup, the objective is to learn the missing physics, $\psi_{\text{missing}}(\bm{F})$ by using the force-displacement data as described before.
3 Proposed approach
In this section, we will present the detailed architecture of the proposed framework. This includes the integration of model and data, the design of the physics-based neural network architecture, and the implementation of correction terms to ensure compliance with the physical constraints outlined in Section. 2.2. Additionally, we will describe the algorithm, provide a flow diagram, discuss the loss function, elaborate on the ensembling and Bayesian inference methods, and detail the training process.
3.1 Architecture design
In a realistic setting, while developing a constitutive model for hyperelastic materials, we generally have access to a low-fidelity phenomenological model and sparse data. In this work, we propose a framework that integrates both sources of information. The fundamental idea is to augment the partially correct phenomenological model (base phenomenological model) with a data-driven deep learning architecture, which can be expressed as:
$\psi(\bm{F})=\psi_{\text{known}}(\bm{F})+\psi_{NN}(\bm{F};\bm{\theta})$ (22)
Here, $\psi_{NN}(\bm{F};\theta)$ represents the neural network component, which is responsible for correcting $\psi_{\text{known}}(\bm{F})$ (the known phenomenological model). The hypothesis is that the neural network will learn the $\psi_{\text{missing}}(F)$ , thereby improving the accuracy of the model. We will exploit the augmented strain energy density function $\psi(\bm{F})$ to derive the stress measures, specifically the first Piola-Kirchhoff stress $\bm{P}(\bm{F})$ and the incremental tangent modulus $\mathbb{C}(\bm{F})$, as described by Eq. (11). Therefore, the primary objective is to accurately determine the function $\psi(\bm{F})$. However, this augmented model $\psi(\bm{F})$ must also satisfy all the physical and thermodynamic constraints outlined in Section.2.2.
3.1.1 Fundamental conditions
The thermodynamic consistency is automatically ensured by constructing the stress measures as the gradient of a potential (strain energy density), as defined in Eq. (11). Additionally, if the strain energy density potential is formulated in terms of the invariants of the right Cauchy-Green deformation tensor $\bm{C}=\bm{F}^{T}\bm{F}$, or equivalently, the Green-Lagrange strain tensor $\bm{E}=(\bm{F}^{T}\bm{F}-\bm{I})/2$, denoted as $\psi(\bm{E}(\bm{F}))$, the resulting Cauchy stress tensor is symmetric and inherently satisfies the condition of objectivity. As shown in Section. 2.3, the known phenomenological model is formulated in terms of invariants. Therefore, we also design the network with these same invariants as input to the network. With this setup, the model automatically satisfies the first three fundamental conditions: thermodynamic consistency, objectivity, and the symmetry of the stress tensor.
3.1.2 Polyconvexity
In order to ensure the condition of polyconvexity of the strain energy density potential, we have adopted Input Convex Neural Network (ICNN) [37] as the neural network architecture. In these Neural Networks, the output is convex with respect to its input arguments. This is ensured by using convex and non-decreasing activation functions and non-negative weights. The Non-negative weights result in convex linear transformations in the hidden layers, as non-negative weighted sums of convex functions remain convex. Additionally, non-decreasing activation functions preserve convexity by ensuring that applying a non-decreasing convex function to a convex input results in a convex output, thereby ensuring the ICNN output is convex in the strain invariants. Further details on the ICNN employed in this work are provided in Section. 3.2.
In addition to the specialized architecture, ensuring polyconvexity also requires that the considered invariants themselves be polyconvex. However, in $\psi(\bm{E}(\bm{F}))$, $\bm{E}$ is not always convex in $\bm{F}$, making it challenging to ensure polyconvexity in $\bm{F}$ using Eq. (16). Although, it is observed that the convexity of $\psi$ in $\bm{E}$ implies local convexity in $\bm{F}$, which ensures local material stability [38, 39]. Note that local convexity is preferable over global convexity because it ensures stability without requiring the strain energy density to remain smooth and bounded as $\det(F)\to 0^{+}$. This flexibility allows the strain energy density to approach infinity, implicitly fulfilling the growth condition necessary for physical realism under extreme compressions. In contrast, global convexity would prevent this by mandating a bounded energy response even in these extreme cases and would require explicitly satisfying the growth condition by adding correction terms [33]. Therefore, we also limit the architecture to local convexity only.
3.1.3 Normalisation conditions
In the undeformed configuration ($\mathbf{F}=\mathbf{I}$), it is essential that both the energy and stress normalization conditions are satisfied. This is achieved by incorporating physics-based terms to ensure the energy density vanishes at zero deformation. To achieve energy normalization, we add $\psi_{energy}$ such that
$\psi(\bm{F})\big{|}_{\bm{F}=\bm{I}}=\left\{\psi_{\text{known}}(\bm{F})+\psi_{NN}(\bm{F};\bm{\theta})\right\}_{\big{|}_{\bm{F}=\bm{I}}}=0$ (23)
The known phenomenological models are constructed to automatically satisfy the normalization condition, i.e., $\psi_{\text{known}}(\mathbf{I})=0$. Consequently, the energy term is defined as:
$\psi_{energy}=-\psi_{NN}(\bm{I};\bm{\theta})$ (24)
Furthermore, the stress correction term $\psi_{\text{stress}}$ is included in the potential to ensure stress vanishes under zero deformation. This term is defined as:
$\psi_{\text{stress}}=\bm{H}:\bm{E}$ (25)
Utilizing Eq. (11), we have:
$\bm{P}(\bm{I})=0=\left\{\frac{\partial\psi_{NN}(\bm{F};\bm{\theta})}{\partial\bm{F}}+\bm{F}\bm{H}\right\}_{\big{|}_{\bm{F}=\bm{I}}}\implies\bm{H}=\frac{\partial\psi_{NN}(\bm{F};\bm{\theta})}{\partial\bm{F}}\bigg{|}_{\bm{F}=\bm{I}}$ (26)
where $\bm{H}$ is a constant symmetric matrix. Here, we exploit the fact that $\psi_{\text{known}}(\bm{F})=0$ for $\bm{F}=\bm{I}$ implies its derivative with respect to $\bm{F}$ at $\bm{F}=\bm{I}$ is also zero. Consequently, we obtain the following expression for the strain energy density:
$\psi(\bm{F};\bm{\theta)}=\psi_{\text{known}}(\bm{F})+\psi_{NN}(\bm{F};\bm{\theta})+\psi_{\text{energy}}+\psi_{\text{stress}}$ (27)
This is referred to as ‘augmented strain energy’ and is used in this work.
3.2 ICNN Architecture
Having discussed the fundamental constraints and a systematic procedure for satisfying the same, we proceed to discussing the architecture of the Input Convex Neural Network (ICNN), which is a key component of the proposed approach. The inputs to the ICNN are modified versions of various invariants of $\bm{C}$, as detailed in Section.2.3. The architecture consists of repeated convex layers, each with 64 neurons in the hidden layer. It incorporates skip connections from the input to the hidden layer after each convex layer. This sequence is repeated three times. Convexity is ensured by maintaining positive weights, achieved by applying the softplus function to the weight matrix prior to its use, as shown below:
$\bm{z}^{l}=\sigma_{1}\left(\underbrace{\sigma_{2}\left(\bm{W}^{l}\right)\bm{z}^{l-1}}_{\text{convex-linear}}+\underbrace{\bm{W}_{s}^{l}\bm{z}^{0}}_{\text{skip-connect}}+\underbrace{\bm{b}^{l}}_{\text{bias}}\right)$ (28)
where $\bm{W}^{l}\in\mathbb{R}^{d^{l}\times d^{l-1}}$, $\bm{W}_{s}^{l}\in\mathbb{R}^{d^{l}\times|\bm{z}^{0}|}$ and $\bm{b}^{l}\in\mathbb{R}^{d^{l}}$ are trainable parameters with $d^{l}$ being the dimension of $l^{th}$ hidden layer, and
$\sigma_{1}(z)=\underbrace{\mathcal{\gamma}_{1}\left(\log(1+e^{z})\right)^{2}}_{\text{squared-softplus}},\quad\text{and}\quad\sigma_{2}(z)=\underbrace{\mathcal{\gamma}_{2}\left(\log(1+e^{z})\right)}_{\text{softplus}}$ (29)
respectively, for all $z\in\mathbb{R}$. Here, $\mathcal{\gamma}_{1}>0$ and $\mathcal{\gamma}_{2}>0$ are hyperparameters. The softplus function guarantees $C^{\infty}$ continuity of the strain energy density and its derivatives, ensuring that stress and tangent are sufficiently smooth and continuous. The function $\sigma_{1}$ is taken as the squared softplus to address the issue of the vanishing second derivative. This squaring not only facilitates the calculation of the tangent of the strain energy density but also prevents vanishing gradients during backpropagation in neural network training. Additionally, skip connections are incorporated to mitigate the issues of vanishing gradients, overfitting, and accuracy degradation (or saturation) as the neural network architecture deepens. At the output of the neural network we get the missing component of the strain energy density.
3.3 Training
As discussed in Section 2.4, we have nodal displacement data represented as $u^{t}(r,j)$, where $t=\{1,2,\ldots,n_{t}\}$ denotes the load step, $r=\{1,2,\ldots,n_{d}\}$ indexes the nodal points, and $j=\{1,2\}$ indicates the direction of displacement. We define the degree of freedom at the $r^{\text{th}}$ node in the $j^{\text{th}}$ direction as $\mathcal{D}_{f}=(r,j)$. This encompasses both $\mathcal{D}_{f}^{\text{free}}$: degrees of freedom within the boundary where displacement is not fixed, and $\mathcal{D}_{f}^{\text{fixed}}$: degrees of freedom at the boundaries where displacement is constrained. Additionally, we have reaction force data $R_{\beta}$ available at selected boundary nodes $\beta=1,2,\ldots,n_{\beta}$.
Degrees of freedom corresponding to these $n_{\beta}$ boundary nodal points constitute $\mathcal{D}_{f}^{fixed}$.
Figure 1: Training Setup: A square plate with a hole in the bottom-left corner is subjected to asymmetric biaxial tension under displacement control. The noisy data collected from full-field displacements and reaction forces is used to train the proposed constitutive model.
Next, we divide the geometry into finite elements and write the displacement as a continuous function of spatial coordinates using shape functions $\{N_{r}(X):r=1,\ldots,n_{d}\}$,
$\bm{u}(\bm{X})=\sum_{r=1}^{n_{d}}N_{r}(\bm{X})\bm{u}^{t}_{r}$ (30)
and $N_{r}:\Omega\rightarrow\mathbb{R}$ is associated with the node at reference coordinates $X=\{X_{r}\in\Omega:r=1,\ldots,n_{d}\}$. The deformation gradient is then given by:
$\bm{F}(\bm{X})=\bm{I}+\sum_{r=1}^{n_{d}}\bm{u}_{r}\otimes\nabla N_{r}(\bm{X})$ (31)
where $\nabla$ denotes the gradient operator with respect to the reference coordinates. Similarly, let’s introduce the test function $\bm{v}$ and describe it in terms of shape functions as:
$\bm{v}(\bm{X})=\sum_{r=1}^{n_{d}}N_{r}(\bm{X})v^{t}_{r,j}\quad\text{with}\quad v^{t}_{r,j}=0\quad\forall(r,j)\in\bigcup_{\beta=1}^{n_{\beta}}\mathcal{D}_{f}^{fixed}$ (32)
From the displacement field $\bm{u}(\bm{X})$, we obtain the deformation gradient following Eq. (31), modified invariants of which are input into the ICNN to yield the strain energy density potential. Using Eq. (11), we derive the stress field. In the absence of stress field data, we use physics-informed loss in form of the linear momentum conservation equation to inform and guide the training process. Note that although we use finite element method in this work, other techniques including finite volume and finite difference can also be used.
In the reference domain $\Omega$, the weak form of the linear momentum balance under quasi-static loading (negligible inertia and body forces) is given by:
[…]
where $\mathbf{v}$ is an admissible test function, sufficiently regular and vanishing on the Dirichlet boundary $\partial\Omega_{u}$, and $\hat{\mathbf{t}}$ represents the surface traction on $\partial\Omega_{t}$. In displacement-controlled testing, $\hat{\mathbf{t}}=0$. The weak form is preferred over the strong form due to its reduced sensitivity to noise in the data. Utilizing Eqs. (32) and (33), we obtain:
[…]
Here, $\mathcal{F}_{r,j}$ is defined as the difference between the external force due to applied tractions on Neumann boundaries and the interior force from the constitutive model. Numerical quadrature over the mesh is used to compute the integrals in Eq. (34). By simplifying Eq. (34) separately for fixed and free degrees of freedom, we obtain:
$\mathcal{F}_{r,j}=0\quad\forall(r,j)\in\mathcal{D}_{f}^{\text{free}}\quad\text{and}\quad\sum_{(r,j)\in\mathcal{D}_{f}^{\text{fixed}}}\mathcal{F}_{r,j}=R_{\beta}\quad\forall\,\beta=1,\ldots,n_{\beta}$ (35)
Although we have used finite element implemented using differentiable finite element simulator, the overall approach can be seamlessly integrated with other numerical techniques such as finite volume method and finite difference method.
To optimize the learnable parameters, we derive the loss function using Eq. (35):
$\mathcal{L}=\sum_{t=1}^{n_{l}}\left[\sum_{(r,j)\in\mathcal{D}_{f}^{\text{free}}}(\mathcal{F}^{t}_{r,j})^{2}+\sum_{(r,j)\in\mathcal{D}_{f}^{\text{fixed}}}(\mathcal{F}^{t}_{r,j}-R_{\beta})^{2}\right]$ (36)
In this optimization problem, although the network is convex with respect to its input, it is non-convex concerning the trainable parameters. Additionally, direct stress or strain energy labels are absent, so we use global reaction force data for training. These factors lead to an optimization problem with multiple solutions (multiple local minima), making the solution highly sensitive to initialization. To address this challenge, we trained multiple ICNNs independently with different random initializations. The model with the minimum loss value among all the trained models is selected and used for prediction. The training algorithm is shown in Algorithm 1.
Input: Point-wise displacement data $U=\{\bm{u}^{t}_{r,j}\},e\in\mathbb{R}^{2},r=1,\ldots,n_{d},t=1,\ldots,n_{l}$
Input: Global reaction forces $R^{\beta}=\{R^{t}_{r,j},\beta=1,\ldots,n_{i},t=1,\ldots,n_{t}\}$
1 Randomly initialize ICNN parameters $\theta$ and the Adam optimizer with a learning rate scheduler;
2 for $e=1,\ldots,epochs$ do
3       $\mathcal{L}\leftarrow 0$ ;
4        // Initialize loss for current epoch
5       for $t=1,\ldots,n_{l}$ do
6             for each element in mesh do
7                   $\psi_{energy}\leftarrow-\psi_{NN}(\bm{F};\bm{\theta})_{\big{|}_{\bm{F}=\bm{I}}}$ ;
8                    // see 24
9                   $\psi_{\text{stress}}\leftarrow\bm{H}:\bm{E}$ where $\bm{H}\leftarrow\frac{\partial\psi_{NN}(\bm{F};\bm{\theta})}{\partial\bm{F}}\bigg{|}_{\bm{F}=\bm{I}}$ ;
10                    // see 25
11                   $\psi(\bm{F};\bm{\theta})\leftarrow\psi_{\text{known}}(\bm{F})+\psi_{NN}(\bm{F};\bm{\theta})+\psi_{energy}+\psi_{\text{stress}}$;
12                    // see 27
13                   $\bm{P}(\bm{F};\bm{\theta})\leftarrow\frac{\partial\psi(\bm{F};\bm{\theta})}{\partial\bm{F}}$
14             end for
15            for $r=1,\ldots,n_{d}$ do
16                   for $j=1,2$ do
17                         Compute force $\mathcal{F}^{t}_{r,j}$; ;
18                          // see (34)
19
20                   end for
21
22             end for
23            for $(r,j)\in\mathcal{D}_{f}^{free}$ do
24                   $\mathcal{L}\leftarrow\mathcal{L}+\left(\mathcal{F}^{t}_{r,j}\right)^{2}$;
25                    // see35 and 35
26
27             end for
28
29       end for
30      for $\beta=1,\ldots,n_{\beta}$ do
31             $r^{\beta,t}=0$;
32             for $(r,j)\in\mathcal{D}_{f}^{fixed}$ do
33                   $r^{\beta,t}\leftarrow r^{\beta,t}+\mathcal{F}^{t}_{r,j}$
34             end for
35            $\mathcal{L}\leftarrow\mathcal{L}+\left(R^{\beta}-r^{\beta,t}\right)^{2}$;
36              // see35 and 35
37
38       end for
39
40 end for
41Compute gradients $\partial\mathcal{L}/\partial\theta$ using automatic differentiation;
42 Update $\theta$ with Adam optimizer using gradients $\partial\mathcal{L}/\partial\theta$;
Output: Trained hybrid model $\psi(\bm{F};\bm{\theta^{*}})$ with optimised parameters $\bm{\theta^{*}}$
Algorithm 1 Training Algorithm for FuCe
After selecting the best-converged network, it is still important to quantify the uncertainty in the predictions. To achieve this, we used approximate Bayesian inference during prediction stage by employing the Monte Carlo dropout method. Details on the same is provided in the next section.
3.4 Approximate Bayesian inference using Monte Carlo dropout
Bayesian approaches offer a powerful framework by providing both predictions and their associated uncertainty. Rather than employing complex Bayesian methods, we utilize Monte Carlo dropout, which is a simpler yet effective approximate Bayesian technique for capturing predictive uncertainty. This approach uses dropout during predictions and runs several forward passes with different dropout masks, that results in producing a distribution of predictions.
Consequently, $N_{m}$ dropout masks are applied to the best-trained model, generating $N_{m}$ (taken as $100$) number of outputs for strain energy density $\hat{\psi}(\bm{F};\bm{\theta})$ and derived stress measure Piola-Kirchhoff stress $\bm{\hat{P}}(\bm{F};\bm{\theta})$ , denoted as $\{\hat{\psi}_{1},\hat{\psi}_{2},\dots,\hat{\psi}_{N_{m}}\}$ and $\{\hat{\bm{P}}_{1},\hat{\bm{P}}_{2},\dots,\hat{\bm{P}}_{N_{m}}\}$. The predicted outputs for a test input $\bm{F}$ during the $n^{\text{th}}$ forward pass of dropout can be expressed as:
$\hat{\psi_{n}}=\psi(\bm{F};\bm{\theta}_{n}^{*},D_{n})\quad\text{and}\quad\hat{\bm{P}_{n}}=\frac{\partial\psi(\bm{F};\bm{\theta}_{n}^{*},D_{n})}{\partial\bm{F}}$ (37)
where $n=\{1,2,\dots,N_{m}\}$ and $\bm{\theta}_{n}^{*}$ represents the parameters of the trained model after applying $n^{\text{th}}$ dropout mask, represented by $D_{n}$, each element of which is a Bernoulli random variable. After obtaining the $N_{m}$ predicted outputs, the expectation of the predicted outputs is estimated as:
$\mathbb{E}[\hat{\psi}]=\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\hat{\psi_{n}}\quad\text{and}\quad\mathbb{E}[\hat{\bm{P}}]=\frac{1}{N_{m}}\sum_{n=1}^{N_{m}}\hat{\bm{P}_{n}}$ (38)
We then calculate the empirical variance of these outputs to estimate the predictive uncertainty of the model as follows:
$\operatorname{Var}(\hat{\psi_{n}})=\frac{1}{N_{m}-1}\sum_{n=1}^{N_{m}}(\hat{\psi_{n}}-\mathbb{E}[\hat{\psi_{n}}])^{2}\quad\text{and}\quad\operatorname{Var}(\hat{\bm{P}})=\frac{1}{N_{m}-1}\sum_{n=1}^{N_{m}}(\hat{\bm{P}_{n}}-\mathbb{E}[\hat{\bm{P}_{n}}])^{2}$ (39)
We quantify the model’s epistemic uncertainty by calculating the empirical variance of these outputs, enhancing the trustworthiness of the model’s predictions, particularly in scenarios with sparse data. This method strengthens the predictive capability of our framework and aligns with the increasing demand for models that offer predictive uncertainty alongside predictions, thereby supporting more informed decision-making. The overall framework proposed in this work is shown in Fig. 2.
Figure 2: The overall framework of the FuCe. Note that for training the proposed approach, differentiable physics simulator is needed that allows backpropagation through the numerical solver.
4 Results and Discussion
In this section, we demonstrate the effectiveness of our approach through various numerical examples. We investigate the effectiveness of the proposed approach in learning the constitutive relation from sparse and noisy data. Additionally, we also showcase its practical utility by fusing the learning hybrid model with finite element solver FeNICS[40]
4.1 Simple Stress States
In this section, we present a comparative evaluation of the proposed data-model fusion framework, the known phenomenological model, and a purely data-driven model without fusion, all benchmarked against the ground truth. The performance of these models is assessed across six stress states: simple shear (SS), pure shear (PS), uniaxial tension (UT), uniaxial compression (UC), biaxial tension (BT), and biaxial compression (BC), with the corresponding deformation gradients provided below:
$\mathbf{F}^{\text{UT}}(\gamma)$ $=\begin{bmatrix}1+\gamma&0\\ 0&1\end{bmatrix};\quad\mathbf{F}^{\text{UC}}(\gamma)=\begin{bmatrix}\frac{1}{1+\gamma}&0\\ 0&1\end{bmatrix},$ (40)
$\mathbf{F}^{\text{BT}}(\gamma)$ $=\begin{bmatrix}1+\gamma&0\\ 0&1+\gamma\end{bmatrix},\quad\mathbf{F}^{\text{BC}}(\gamma)=\begin{bmatrix}\frac{1}{1+\gamma}&0\\ 0&\frac{1}{1+\gamma}\end{bmatrix},$
$\mathbf{F}^{\text{SS}}(\gamma)$ $=\begin{bmatrix}1&\gamma\\ 0&1\end{bmatrix},\quad\mathbf{F}^{\text{PS}}(\gamma)=\begin{bmatrix}1+\gamma&0\\ 0&\frac{1}{1+\gamma}\end{bmatrix}.$
where $\gamma$ represents the loading parameter. Note that these deformation paths are used only during prediction for evaluating the model. During training, a single model is trained based on a complex generalized stress states, rather than on individual deformation paths. We have utilized FEM-emulated digital image correlation data from [33]. To emulate a realistic scenario, the dataset is corrupted with Gaussian white noise. Further, to illustrate the capability of the proposed approach in handling small and sparse dataset, we only utilized 50% of the load steps. We evaluate the performance of the proposed constitutive framework using three different elastic solids, as detailed in Section 4.1.1, Section 4.1.2, and Section 4.1.3. To assess the model’s extrapolative capabilities, $\gamma$ is extended beyond its training range of $\gamma\in[0,1]$. Furthermore, when employing the approximate Bayesian method to predict the mean output and confidence interval, a dropout rate of $0.5$ is used during Monte Carlo dropout.
4.1.1 Isihara
In this case, we have chosen the Isihara hyperelastic model as our ground truth. Originally formulated by Isihara et al. [35], this model is renowned for its high accuracy in predicting the response of hyperelastic materials. The enhanced accuracy is attributed to the incorporation of the second invariant in the formulation of the strain energy function, as presented in Eq. (18). We considered the Neo-Hookean model (Eq. (17)) as the known phenomenological model $\psi_{\text{known}}$ and trained a single deep learning model $\psi_{NN}$ on generalized state of stress (with only four load steps) to learn the discrepancies. To enable training the model directly from displacement measurements, we developed a differentiable physics solver that allows backpropagation through the solver. This allows us to train the model directly from the displacement data. The model is trained based for $\gamma\in[0,1]$; however, during evaluation, we considered $\gamma\in[0,3]$ to investigate the extrapolation performance of the model. Additionally, the trained model is evaluated in six different state of stress mentioned earlier.
Figures 3(a) and 3(b) illustrate the variation in strain energy density and the Piola-Kirchhoff stress tensor, respectively, as functions of the loading parameter $\gamma$ for six deformation path mentioned above. The proposed framework consistently outperforms the other models across all deformation paths, except for BT. One potential solution towards improving the prediction for BT is to fine-tune the model for BT data; however, the same is not explored in the work. We note that purely data-driven models also perform well for $\gamma\in[0,1]$; however, it starts deviating as we increase $\gamma$ beyond 1. In contrast, the proposed framework, leveraging data-model fusion, sustains its accuracy even in regions extending three times beyond the training window. This clearly demonstrates the advantage of data-physics fusion over purely data-driven model.
Figure 3: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the the Isihara model serves as the ground truth.
4.1.2 Arruda-Boyce
Figure 4: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the the Arruda-Boyce model serves as the ground truth.
In this case, we have selected the Arruda-Boyce hyperelastic model as the ground truth. Developed by Arruda and Boyce [36], this model is also referred to as the 8-chain model due to its distinctive network structure, which comprises eight chains extending from the center of a cubic unit cell to its corners. The Neo-Hookean model (Eq. (17)) is again employed as the known phenomenological model ($\psi_{\text{known}}$), and a deep learning model $\psi_{NN}$ is trained on a generalized state of stress, utilising data from only first six load steps, to learn the discrepancies. Again, we note that the model is trained based on displacement data by exploiting the differentiable solver that allows backpropagation through the solver. To assess the model’s extrapolation capabilities, we extended the loading parameter to $\gamma\in[0,3]$ and evaluated the trained model across six different stress states as before.
Figures 4(a) and 4(b) show the strain energy density and Piola-Kirchhoff stress tensor as functions of the loading parameter $\gamma$ across six deformation paths, comparing the proposed framework’s performance with other models. The framework outperforms the other models, particularly in the extrapolated region where $\gamma>1$. For the deformation paths of Pure Shear (PS) and Simple Shear (SS), where both the data-driven model and the known Neo-Hookean model deviate significantly from the ground truth in the extrapolative region, the proposed model closely aligns with the true values, nearly encompassing them. Notably, in the case of the uni-axial tension (UT), the proposed model follows its data-driven counterpart, which coincides with the ground truth. Similarly, for the bi-axial tension (BT), it aligns with its known phenomenological counterpart. This behavior highlights the effectiveness of data-model fusion, leveraging the strengths of both approaches. It is also noteworthy that for the UT and BT deformation paths, the confidence intervals are narrow and are not discernible in the figures. This can be attributed to small variation in outputs for these deformation paths across all dropout masks, resulting in very small confidence intervals.
4.1.3 Anisotropic solid with one fiber family oriented at $45^{o}$
Figure 5: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model, along with the known fiber direction $\alpha$, while the anisotropic model with a single fiber family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
Figure 6: Comparison of predicted outputs: (a) Strain energy density $\psi(F(\gamma))$ and (b) First Piola–Kirchhoff stress $\bm{P}(\bm{F}(\gamma))$, produced by the proposed constitutive model, against a data-driven model without fusion and a known phenomenological model, relative to the ground truth across six distinct stress states. The Neo-Hookean model is utilized as the known phenomenological model while the anisotropic model with a single fiber family aligned at $\alpha=45^{\circ}$ serves as the ground truth.
In this case, we selected an anisotropic material with a single fiber family oriented at 45 degrees as our ground truth. The Neo-Hookean model (Eq. (17)) is again used as the known phenomenological model, $\psi_{\text{known}}$, while a deep learning model, $\psi_{NN}$, is trained on generalized stress states using data from only the first five load steps to capture the discrepancies. A key aspect of this anisotropic case is the fiber orientation, denoted as $\alpha$. Two scenarios were investigated: in the first, the fiber direction $\alpha$ was predefined; in the second, it was treated as a learnable parameter, as described below:
$\psi=\psi(\bm{F};\bm{\theta}_{\alpha})=\psi_{\text{known}}(\bm{F})+\psi_{\text{NN}}(\bm{F};\bm{\theta}_{\alpha})+\psi_{\text{energy}}+\psi_{\text{stress}}$
where $\bm{\theta}_{\alpha}$ represents the set of learnable parameters, including $\alpha$. It is introduced as an input to the neural network with the invariant $(\tilde{I}_{\alpha}-1)^{2}$ where $\tilde{I}_{\alpha}=J^{-2/3}(\mathbf{a}\cdot\bm{C}\mathbf{a})$ and $\mathbf{a}=(\cos\alpha,\sin\alpha,0)^{T}$ as mentioned earlier. The network is trained to optimize both $\alpha$ and the weights and biases $\theta$ using gradient-based methods and gradients with respect to both $\alpha$ ($\frac{\partial\mathcal{L}}{\partial\alpha}$) and $\bm{\theta}$( $\frac{\partial\mathcal{L}}{\partial\bm{\theta}}$) are computed to minimize the loss function during training. Furthermore, the loading parameter $\gamma$ was extended beyond its training range up to $\gamma\in[0,1.7]$ in both scenarios to assess the framework’s extrapolation capabilities.
In the first scenario, Figures 5(a) and 5(b) illustrate the good alignment of the proposed approach with the ground truth, clearly outperforming other models across all stress states, particularly in the extrapolated regions.
In the second scenario, Figures 6(a) and 6(b) highlight the performance of the proposed framework across all the stress states, demonstrating a close alignment with the ground truth in both strain energy density and Piola-Kirchhoff stress plots. In this case, accurate learning of fibre direction $\alpha$ is crucial for the accuracy of the model; therefore, we incorporated polar plots alongside the $\psi(\gamma)$ and $P(\gamma)$ plots to compare the learned fiber direction $\alpha$ from the data-driven and proposed approaches relative to the ground truth. It is noteworthy that the Neo-Hookean phenomenological model, lacking directional information, does not contribute to the polar plot. The results indicate that the learned $\alpha$ closely matches the true $\alpha$, as evidenced in the polar plot and the proposed framework significantly outperforms the other models in the $\psi(\gamma)$ and $P(\gamma)$ plots, particularly in the extrapolated region where $\gamma>1$. Additionally, the confidence interval in most stress states is too narrow to be noticeable visually. This is likely due to very small variation in the prediction outputs for these stress states across the dropout masks, resulting in very small confidence intervals.
4.2 FEM Analysis
(a) Mesh-1
(b) Mesh-2
(c) Mesh-3
(d) Ground Truth-1
(e) Ground Truth-2
(f) Ground Truth-3
(g) Proposed-1
(h) Proposed-2
(i) Proposed-3
Figure 7: FEM simulation results, using the proposed constitutive model with Neo-Hookean as the known phenomenological model and Isihara model as the ground truth.
[…]
The domain for each geometry was discretized using a mesh with 64 elements in each direction. For the cases with holes, additional mesh refinement was performed around the holes to better capture the stress concentrations. The FEM analysis was performed using the FEniCS package, incorporating the proposed framework as the constitutive model. The results were benchmarked against those derived from the exact Isihara phenomenological model, as illustrated in Fig. 7. The results indicate that the FEM analysis using the proposed framework matches well with the phenomenological model, demonstrating the effectiveness of the proposed approach in handling complex geometries and loading conditions.
5 Conclusion
In this paper, we present a novel hybrid framework, "Fusion-based Constitutive model (FuCe)", designed for constitutive modeling with a focus on hyperelastic materials. Our approach seamlessly integrates existing phenomenological constitutive models with Input Convex Neural Networks (ICNNs), enabling the deep learning model to capture the discrepancy between the traditional phenomenological model and the actual material behavior as observed in the data. This method not only corrects the phenomenological model based on data but also preserves the model’s interpretability and generalization capabilities.
We applied the proposed framework to two isotropic and one anisotropic materials material. The framework was rigorously evaluated under six different stress states and further validated through finite element method (FEM) simulations for Isihara material to assess its practical applicability. The key findings from this study are summarized below:
  • The proposed framework demonstrates good accuracy in all the example problems, along with the FEM simulation. The accuracy is maintained even for extrapolations in terms of loading parameter. This showcases the robustness, generalization capacity of the proposed framework.
  • The proposed framework satisfies all essential constitutive requirements, including material objectivity, stress symmetry, local polyconvexity, and normalization conditions. Some of these are inherently achieved by construction, while others are ensured by incorporating physics-based corrective terms. This approach guarantees stress-free reference states and aligns the model closely with the behavior of real-world materials
  • The proposed model is trained using practically feasible data, consisting of sparse and noisy full-field displacement measurements and global reaction force data. This training process leverages a differentiable solver, enabling backpropagation through the solver. In the absence of stress labels, the model’s learning is guided by a physics-informed loss function based on the conservation of linear momentum, effectively compensating for the lack of direct stress measurements.
  • The proposed framework also quantifies predictive uncertainty by incorporating Monte Carlo dropout, a simple yet effective approximate Bayesian method. This approach enhances the dependability of the model’s predictions and supports more informed decision-making.
In summary, the proposed FuCe framework successfully integrates classical phenomenological models with emerging deep learning approaches for constitutive modeling of complex materials, such as hyperelastic materials, as investigated in this study. It demonstrates strong alignment with ground truth data, even in extrapolative regions, and is satisfactorily applicable within FEM simulations, producing accurate results. However, some deviations were observed, particularly for the biaxial tensile stress-state. One potential solution is to fine tune FuCe for biaxial data. Additionally, we have only focused on hyperelastic material model; other constitutive relations have not been investigated in this work. Future work will focus on addressing these challenge.
Acknowledgements
SC acknowledges the financial support received from Anusandhan National Research Foundation (ANRF) via grant no. CRG/2023/007667. SC and RN acknowledges the financial support received from the Ministry of Port and Shipping via letter no. ST-14011/74/MT (356529).
Code availability
Upon acceptance, all the source codes to reproduce the results in this study will be made available on request
Competing interests
The authors declare no competing interests.